Circular No. 7897

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2002 J1 AND C/2002 J2 (SOHO) Further to IAUC 7886, D. Hammer reports two Kreutz-sungrazing

comets found by R. Kracht on SOHO C2 images (cf. MPECs cited below).

Comet  2002 UT  α_{2000}  δ_{2000}  MPEC
C/2002 J1 May 1.726  2^{h}40.^{m}6  +14^{o}01′  2002-J09
C/2002 J2  3.476  2 48.0  +14 35  2002-J29

XTE J1908+094 S. Chaty, Open University; and R. P. Mignani, European Southern

Observatory (ESO), Garching, report on the detection in the near-infrared of a likely counterpart to the blackhole candidate XTE J1908+094 (IAUC 7856). Target-of-opportunity observations performed by L. Vanzi (ESO) with the ESO New Technology Telescope on Apr. 25 and 29 show an infrared source that is coincident (within an uncertainty of ± 0′′.17 in each coordinate) with the position of the radio counterpart (IAUC 7874). The magnitudes of the source on Apr. 25 were J = 18.7 ± 0.3, H = 17.2 ± 0.3, K_{s} = 16.4 ± 0.2. Although the source was also visible on Apr. 29, the data do not allow an accurate determination of the magnitude and therefore prevent any variability measurement. Taking the value of column density given on IAUC 7856, and assuming that the source is located near the Galactic center, absolute magnitudes of J = 0.4, H = 0.3, and K = 0.3 are derived. Chaty and Mignani add: “If the source is indeed the actual counterpart of XTE J1908+094, its infrared flux could be significantly contaminated by the accretion disk. In this case, the derived upper limits show that the source is a low-mass system with a companion star of spectral type later than A. Further photometric observations are encouraged to search for variability of this source.”

XTE J0929–314 D. K. Galloway, E. H. Morgan, R. A. Remillard, and D. Chakrabarty,

Massachusetts Institute of Technology, write: “Observations of the recently discovered 185-Hz accreting millisecond pulsar XTE J0929–314 (cf. IAUC 7888, 7889, 7893, 7895) with RXTE/PCA on May 9.5 UT indicate that the pulsar is still active at a flux level of ∼ 20 mCrab (2–10 keV), down from 30 mCrab on May 2. Orbital Doppler shifts of the pulse frequency were clearly detected. Further RXTE observations are scheduled over the next few days. Multiwavelength follow-up observations are strongly encouraged.”

2002 May 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green